Molarity and molality of a solution of a liquid (mol wt.50) in aqueous solution is 9 and 10 respectively . What is the density of solution??

Grade:12th pass

1 Answers

Vikas TU
7 years ago
Molarity = moles of solute/volume of soln. in ltres.
      9 mol/litre     = (w/50)/(m/d)...................................(1)
10 molality means 10 moles are present in 1 kg water.
molar mass = 500 gm is present in 100 gm water. 
mass of solution = 1500 gm.
Volume = 9M/10mol and volume  = 1.1111 l
density comes to =  1500/1111 = 1.350 g/l